If a physical inventory was taken in January and the actual inventory is $209,400, is there a shortage or overage and by how much in dollars? Net sales for the year is $153,800.

  • January BOM                        $215,480
  • January Sales                     $26,675
  • January Markdowns         $24,485
  • January Receipts             $48,500
Answer #1

Solution:

EOM Stock = [BOM - Net Sales - Net Markdowns + Net Receipts]

EOM Stock = 215,480 - 26675 - 24,485 + 48,500 = 212,820

Actual inventory = 209,400

Shortage = 212,820 - 209,400 = 3,420

Percentage : 3,420 / 153,800 * 100 = 2.22%